# concert_supp

### v2

@@ 3 Variational Bayes Implementation
The proposed Bayesian transfer learning can be carried out using standard Gibbs sampling from full conditional posterior distributions. However, here we adopt a scalable approach based on variational Bayes (Ray et al., 2020; Ray and Szabó, 2022). Variational Bayes (VB) serves as a popular tool to improve Bayesian methods on the computational cost of MCMC (Blei et al., 2017). The thrust of VB is to approximate the joint posterior distribution by the closest distribution within some tractable distribution family. VB is commonly used to assist stochastic search variable selection for spike-and-slab prior (Ray et al., 2020; Menacher et al., 2023), where its performance is guaranteed both empirically and theoretically (Carbonetto and Stephens, 2012; Ray and Szabó, 2022). In this section, we derive the VB algorithm for transfer learning.
We denote the likelihood functions of the target data and the source data as $L({\cal D}^{(0)})$ and $L({\cal D}^{(k)})$ for $k=1,\ldots,K$, respectively. Let $\theta^{(0)}$ and $\theta^{(k)}$ for $k=1,\ldots,K$ denote all parameters in the target and source models and $\Pi(\theta^{(0)})$ and $\Pi(\theta^{(k)})$ be the corresponding prior distributions. The joint prior distribution is thus given by $\Pi(\theta)=\bigotimes^{K}_{k=0}\Pi(\theta^{(k)})$. Let $\cal{D}$ denote all data. By Bayes’ formula the joint posterior density can be expressed as
$d\Pi(\theta\mid{\cal D})\propto L({\cal D}^{(0)})\prod^{K}_{k=1}L({\cal D}^{(k)})d\Pi(\theta).$ (2)
We propose a hierarchical mean-field variational family $\cal Q$ where $\beta^{(0)}_{j}$’s follow independent spike-and-slab distributions for $j=1,\ldots,p$:
$\beta^{(0)}_{j}\mid Z_{j}\sim Z_{j}N(\mu^{(0)}_{j},(\sigma^{(0)}_{j})^{2})+(1-Z_{j})\delta_{0},\quad Z_{j}\sim\operatorname{Bernoulli}(\gamma^{(0)}_{j}),$
and $\beta^{(k)}_{j}$’s are independent from each other while dependent on $\beta^{(0)}_{j}$ for $k=1,\ldots,K$:
@@
We further denote the variational distribution of other parameters belonging to each specific model as $Q(\phi^{(k)})$ for $k=0,\ldots,K$. Then $\cal Q$ can be represented as
$\mathcal{Q}=\left\{Q(\theta)=\bigotimes_{k=0}^{K}\left[\bigotimes_{j=1}^{p}\left(\gamma^{(k)}_{j}N(\mu^{(k)}_{j},(\sigma^{(k)}_{j})^{2})+(1-\gamma^{(k)}_{j})\Delta^{(k)}_{j}\right)\times Q(\phi^{(k)})\right]\right\},$ (3)
where $\Delta^{(k)}_{j}$ denotes the corresponding Dirac mass distributions. The variational posterior $Q^{*}$ is the element of $\cal Q$ which minimizes the KL divergence to the exact posterior distribution, or equivalently maximizes the evidence lower bound (ELBO):
$Q^{*}=\operatorname{argmax}_{Q\in{\cal Q}}\operatorname{ELBO}\equiv\operatorname{argmax}_{Q\in{\cal Q}}\mathbb{E}_{Q}\left[\log\frac{d\Pi(\theta\mid{\cal D})}{dQ(\theta)}\right].$ (4)
While the exact posterior distribution contains a sum of $2^{(K+1)p}$ models, the VB family $\cal Q$ replaces them with only $(K+1)p$ marginal models, drastically reducing the computational burden in the high-dimensional case.
@@ 3.1 Example: Linear Regression
Under the linear regression setting, the target and source models satisfy $y_{i}^{(k)}=x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}+\varepsilon_{i}^{(k)}$, where $\varepsilon_{i}^{(k)}$’s are i.i.d. random noises from $N(0,(\sigma^{(k)}_{y})^{2})$ for $k=0,\ldots,K$, respectively. We choose a commonly used inverse-Gamma prior distribution $\operatorname{I\Gamma}(a_{0},b_{0})$ on the additional parameters $\{(\sigma^{(k)}_{y})^{2}\}^{K}_{k=0}$, and assume the prior slab variances of the target and sources contain $(\sigma^{(k)}_{y})^{2}$ in addition to $\eta$ and $\tau_{k}$ for conjugacy, respectively. The hyperparameters in the weakly informative inverse-Gamma prior distributions are set as $a_{0}=2$ and $b_{0}=1$ in our numerical studies. For the VB implementation, the variational distributions for $(\sigma^{(k)}_{y})^{2}$ in (3) are assumed to be inverse-Gamma distributions $\operatorname{I\Gamma}(a^{(k)},b^{(k)})$ for $k=0,\ldots,K$.
Since the assigned prior distributions are conditionally conjugate in the linear regression scenario, coordinate ascent variational inference (CAVI) (Bishop, 2006) can be used to solve (4), which iteratively optimizes each factorized term while keeping the others fixed. We defer explicit updates and implementation details to Supplementary Material Section LABEL:supp:algo.
@@ 3.2 Example: Logistic Regression
Unlike linear regression models, the logistic regression models do not enjoy direct conjugacy between likelihood and spike-and-slab priors, which makes MCMC and VB schemes intractable. One way to overcome this difficulty is to adopt Pólya-Gamma (PG) data augmentation (Polson et al., 2013). We introduce a latent PG variable $\omega^{(k)}_{i}$ for each binary response $y^{(k)}_{i}$ and therefore the augmented likelihood $L({\cal D}\mid\omega)$ is given by
@@
where $\omega=(\omega^{(0)}_{1},\ldots,\omega^{(K)}_{n_{K}})$ with each $\omega^{(k)}_{i}$ following the PG distribution $\operatorname{PG}(1,0)$ for $i=1,\ldots,n_{k}$ and $k=0,\ldots,K$. The corresponding variational distributions of $\omega^{(k)}_{i}$’s are assumed to be $\operatorname{PG}(1,c^{(k)}_{i})$. Data augmentation guarantees conjugacy so that CAVI can be utilized to solve (4). A more detailed derivation can be found in Supplementary Material Section LABEL:supp:algo.
@@ 4 Theoretical Results
In this section, we study the theoretical properties of CONCERT under the linear regression setting. Without loss of generality, we assume that $\sigma_{y}^{(0)}=\cdots=\sigma_{y}^{(K)}$ is known in our theoretical analysis. All the proofs can be found in Supplementary Material Section LABEL:app:theorem.
@@
We first introduce some notations. Let $[K]=\{1,\ldots,K\}$ and $N=\sum_{k=0}^{K}n_{k}$. Let $S=\{j\in[p]:\beta_{j}^{(0)}\neq 0\}$ be the target signal set. We denote $T_{k}=\{j\in[p]:\delta_{j}^{(k)}\neq 0\}$ to be the non-transferable index set for source $k$, and $T=\{T_{k}\}_{k=1}^{K}$. We consider the models with $|S|\leq L_{s}$ and $|T_{k}|\leq L_{t}$ for some large numbers $L_{s}$ and $L_{t}$. Let $\{\boldsymbol{\beta}^{(0)*}$, $\{\boldsymbol{\delta}^{(k)*}\}_{k=1}^{K}\}$ be the true parameters. We define the true signal and non-transferable index sets as
$S^{*}=\{j\in[p]:|\beta_{j}^{(0)*}|\geq\zeta_{j}\}\quad\text{and}\quad T^{*}_{k}=\{j\in[p]:\delta_{j}^{(k)*}\neq 0\},$ (5)
respectively, with $s^{*}=|S^{*}|$ and $t_{k}^{*}=|T_{k}^{*}|$ denoting the true target model size and true non-transferable model size of source $k$ for $k\in[K]$.
@@
For a matrix $A$ and an index set $D$, let the sub-matrix of $A$ with columns in $D$ as $A_{D}$. Denote the design matrix on the $k$th dataset as $X^{(k)}$. Then $\tilde{P}_{T_{k}}=X^{(k)}_{T_{k}}(X^{(k)\top}_{T_{k}}X^{(k)}_{T_{k}})^{-1}X^{(k)\top}_{T_{k}}$ and $\tilde{Q}_{T_{k}}=I_{n_{k}}-\tilde{P}_{T_{k}}$ are the projection matrices on the column space of $X_{T_{k}}^{(k)}$ and its complementary space, respectively. Let $\tilde{Q}_{T_{0}}=I_{n_{0}}$ and $\tilde{Q}$ be the diagonal block matrix of $\{\tilde{Q}_{T_{k}}\}_{k=0}^{K}$. We define the transformed design matrix as $\tilde{X}^{(k)}(T_{k})=\tilde{Q}_{T_{k}}^{1/2}X^{(k)}$ and the stacked transformed design matrix as $\tilde{\boldsymbol{X}}(T)=({X}^{(0)\top},\tilde{X}^{(1)}(T_{1})^{\top},\ldots,\tilde{X}^{(K)}(T_{K})^{\top})^{\top}\in\mathbb{R}^{N\times p}$. We further define $\Phi_{S,T}=\tilde{\boldsymbol{X}}_{S}(T)(\tilde{\boldsymbol{X}}_{S}(T)^{\top}\tilde{\boldsymbol{X}}_{S}(T))^{-1}\tilde{\boldsymbol{X}}_{S}(T)^{\top}$.
@@
Given a model $\{S,T\}$, we define the covariate-$j$-specific transferable set as $H^{(j)}_{S,T}=\{k\in[K]:j\in S\setminus T_{k}\}\cup\{0\}$. Denote $H_{S,T}=\cup_{j\in S}H_{S,T}^{(j)}$ as the index set of all least-transferable sources given $\{S,T\}$, indicating that all sources in $H_{S,T}$ have at least one covariate similar to the target. We further define $n^{(j)}_{S,T}=\sum_{k\in H^{(j)}_{S,T}}n_{k}$ to be the useful sample size for covariate $j$ and $n_{S,T}=\sum_{k\in H_{S,T}}n_{k}$ as the sample size of all least-transferable sources given model $\{S,T\}$. Similarly to $\tilde{\boldsymbol{X}}(T)$ defined above, we use $\tilde{\boldsymbol{X}}_{H_{S,T}}$ to denote the stacked design matrix consisting of all the least-transferable sources in $H_{S,T}$.
4.1 Oracle theoretical results with known $T^{*}$
@@
[…]
Assumptions 1(i)-(ii) require that the order of the total signal strength of each dataset is no more than $n_{k}\log p$ and the non-influential signal strength in the transformed dataset is not large (Yang et al., 2016). Assumption 1(iii) requires that the transferable sample size along each dimension is comparable to the total sample size, excluding true models where only a small portion of sources are useful along each dimension.
Remark 2.
The definition of $T_{k}^{*}$ in (5) requires that $\|\boldsymbol{\delta}^{(k)}_{T_{k}^{*\mathrm{c}}}\|_{2}=0$. With more precise analysis, we allow for approximately similar sources with $\|\boldsymbol{\delta}^{(k)}_{T_{k}^{*\mathrm{c}}}\|^{2}_{2}\lesssim(\sigma^{(0)}_{y})^{2}\log p/n_{k}$, which aligns with the spirit of Assumption 1(ii). We assume that they are zero for simplicity.
@@
(ii) (Restricted eigenvalue) The gram matrix of the considered model space satisfies that $\min_{\{S:|S|\leq L_{s}\}}\lambda_{\min}({\tilde{\boldsymbol{X}}_{H_{S},S}^{\top}\tilde{\boldsymbol{X}}_{H_{S},S}}/{n_{S}})\geq\lambda_{l}$ for some $\lambda_{l}\in(0,1)$, $\lambda_{\operatorname{min}}(X^{(0)\top}_{S^{*}}X^{(0)}_{S^{*}})\geq n_{0}\lambda_{l}$, and $\lambda_{\operatorname{min}}(X^{(k)\top}_{T_{k}^{*}}X^{(k)}_{T_{k}^{*}})\geq n_{k}\lambda_{l}$ for all $k\in[K]$ with $|T_{k}^{*}|\geq 1$;
@@
(iv) (Weak correlation) For all $k\in[K]$, and $j\in[p]$, the transformed covariates satisfy that $\|\tilde{X}_{j}^{(k)}\|^{2}_{2}\geq d_{1}n_{k}$ and $\tilde{\boldsymbol{X}}_{j}^{\top}(I-\Phi_{S})\tilde{\boldsymbol{X}}_{j}\geq d_{1}d_{2}n^{(j)}_{S}$ for all $S$ satisfying $|S|\leq L_{s}$, with some $d_{1},d_{2}\in(0,1)$.
Assumption 2 imposes conditions on the projected design, which are parallel to the assumptions for single-task high-dimensional regression in the Bayesian literature (Narisetty and He, 2014; Yang et al., 2016). Assumption 2(ii) allows the model size $S$ to grow slowly compared to $p$ and is satisfied when $L_{s}\leq\min\{p,n/\log p\}$ and Assumption 2(iv) holds. Assumptions 2(iii) and 2(iv) are mild and can hold with $C_{3}=4\lambda_{l}^{-1}$ and $d_{1}=d_{2}=1$ under orthogonal design.
@@
Let $\underline{\tau}=\min_{k}|{\tau}_{k}|$. Assume that the prior satisfies (i) $q_{0}\leq p^{-c}$ with $c>2+{3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/2+C_{2}+C_{3}$; (ii) $\max_{j\in S^{*}}n^{(j)}_{S^{*}}\eta^{2}\leq Cp^{2r}$ for some $r>0$; (iii) $n_{k}\tau_{k}^{2}\leq Cp^{2r_{k}}$ for some $r_{k}>0$ for $k\in[K]$.
Assumption 3 is on the choice of prior hyperparameters, which is similarly assumed in Narisetty and He (2014) and Narisetty et al. (2019).
@@
$\zeta_{j}^{2}\geq C_{\beta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n^{(j)}_{S^{*},T^{*}}d_{1}d_{2}},\text{ for all }j\in S^{*},$
with constant $C_{\beta}=8({3}C_{1}{\lambda_{l}}^{-1}(\eta^{-2}+{\underline{\tau}}^{-2})/2+r+c+2)+4(C_{2}+C_{3})$.
Assumption 4 is parallel to the $\beta$-min condition in the high-dimensional literature for variable selection consistency (Bühlmann and Van De Geer, 2011), which requires the signal strength on the significant set to have a large enough magnitude. Note that the minimal signal strength restriction here is covariate-specific, depending on how many transferable sources there are along with the specific dimension.
@@
Suppose $25\max\{1,s^{*}\}\log p\leq\min_{\{S:|S|\leq L_{s}\}}n_{S}$. Under Assumptions 1- 4, the posterior of the true model $S^{*}$ satisfies that,

### v3

@@ 3 Variational Bayes Implementation
The proposed Bayesian transfer learning can be carried out using MCMC sampling from full conditional posterior distributions. However, here we adopt a scalable approach based on variational Bayes (VB), which serves as a popular tool to alleviate the computational burden of MCMC in Bayesian inference (Jaakkola1997; blei2017review). The thrust of VB is to approximate the joint posterior distribution by the closest distribution within some tractable distribution family. VB is commonly used to assist stochastic search variable selection for spike-and-slab prior (Menacher2023vb), where its performance is guaranteed both empirically and theoretically (Carbonetto2012vb; ray2022vb). In this section, we derive the VB algorithm for transfer learning.
We denote the likelihood of the target data and the source data as $L({\cal D}^{(0)})$ and $L({\cal D}^{(k)})$ for $k=1,\ldots,K$, respectively. Let $\boldsymbol{\theta}^{(0)}$ and $\boldsymbol{\theta}^{(k)}$ for $k=1,\ldots,K$ denote all parameters in the target and source models, with corresponding priors $\Pi(\boldsymbol{\theta}^{(0)})$ and $\Pi(\boldsymbol{\theta}^{(k)})$. The joint prior distribution is then given by $\Pi({\boldsymbol{\theta}})=\bigotimes^{K}_{k=0}\Pi(\boldsymbol{\theta}^{(k)})$ with $\boldsymbol{\theta}=(\boldsymbol{\theta}^{(k)})^{K}_{k=0}$. Let $\cal{D}$ denote all data. By Bayes’ formula, the joint posterior density can be expressed as
$d\Pi(\boldsymbol{\theta}\mid{\cal D})\propto L({\cal D}^{(0)})\prod^{K}_{k=1}L({\cal D}^{(k)})d\Pi(\boldsymbol{\theta}).$ (2)
To preserve the spike‐and‐slab structure assumed in the priors, we propose a hierarchical mean-field variational family $\cal Q$, where the source coefficients remain conditional on the target coefficients, extending the framework in ray2022vb to our transfer learning setting. Specifically, in the target model, the variational distributions for $\beta^{(0)}_{j}$’s follow independent spike-and-slab distributions for $j=1,\ldots,p$:
$\beta^{(0)}_{j}\mid Z_{j}\sim Z_{j}N(\mu^{(0)}_{j},(\sigma^{(0)}_{j})^{2})+(1-Z_{j})\delta_{0},\quad Z_{j}\sim\operatorname{Bernoulli}(\gamma^{(0)}_{j})$.
In addition, for each source $k=1,\ldots,K$, the variational distributions of $\beta^{(k)}_{j}$’s can be represented in a conditional spike‐and‐slab structure as:
@@
For other parameters $\phi^{(k)}$ specified by the specific model, we denote the variational distributions as $Q(\phi^{(k)})$, for $k=0,\ldots,K$. For example, in the linear case, $\phi^{(k)}$ represents the noise variance $(\sigma^{(k)}_{y})^{2}$, while in the logistic case, $\phi^{(k)}$ comprises the augmented $\{\omega^{(k)}_{i}\}^{n_{k}}_{i=1}$, as detailed in Section 3.2. Combining these variational distributions, the variational family $\cal Q$ can be written in a compact and unified form for the GLM transfer learning setting as
$\mathcal{Q}=\left\{Q(\boldsymbol{\theta})=\bigotimes_{k=0}^{K}\left[\bigotimes_{j=1}^{p}\left(\gamma^{(k)}_{j}N(\mu^{(k)}_{j},(\sigma^{(k)}_{j})^{2})+(1-\gamma^{(k)}_{j})\Delta^{(k)}_{j}\right)\times Q(\phi^{(k)})\right]\right\},$ (3)
where $\Delta^{(k)}_{j}$ denotes the corresponding Dirac mass distributions with $\Delta^{(0)}_{j}=\delta_{0}(\beta^{(0)}_{j})$ and $\Delta^{(k)}_{j}=\delta_{\beta^{(0)}_{j}}(\beta^{(k)}_{j})$. This factorization allows the variational family to retain the variable selection and similarity selection capabilities while ensuring tractable optimization. The variational posterior $Q^{*}$ is defined as the element in $\cal Q$ that minimizes the KL divergence to the exact posterior, or equivalently maximizes the evidence lower bound:
$Q^{*}=\operatorname{argmax}_{Q\in{\cal Q}}\mathbb{E}_{Q}\left[\log\frac{d\Pi(\boldsymbol{\theta}\mid{\cal D})}{dQ(\boldsymbol{\theta)}}\right].$ (4)
Based on $Q^{*}$, variable selection in the target model is determined by the variational posterior inclusion probabilities $\gamma^{(0)}_{j}$, where a covariate is selected if its posterior inclusion probability exceeds a prespecified threshold. Parameter estimation is given by the variational posterior mean. Specifically, in the spike-and-slab framework, the estimated values of $\beta^{(0)}_{j}$ are given by $\mathbb{E}_{Q^{*}}[\beta^{(0)}_{j}]=\gamma^{(0)}_{j}\mu^{(0)}_{j}$ for $j=1,\ldots,p$.
Next, we present the primary setup of VB for linear and logistic regression, and defer explicit updates and implementation details to Section C of the Supplementary Material.
@@ 3.1 Example: Linear Regression
Under the linear regression setting, the target and source models satisfy $y_{i}^{(k)}=x^{(k)\top}_{i}\boldsymbol{\beta}^{(k)}+\varepsilon_{i}^{(k)}$, where $\varepsilon_{i}^{(k)}$’s are i.i.d. random noises from $N(0,(\sigma^{(k)}_{y})^{2})$ for $k=0,\ldots,K$, respectively. We choose a commonly used inverse-Gamma prior distribution $\operatorname{I\Gamma}(a_{0},b_{0})$ on the additional parameters $\{(\sigma^{(k)}_{y})^{2}\}^{K}_{k=0}$, and assume that the prior slab variances of the target and sources contain $(\sigma^{(k)}_{y})^{2}$ in addition to $\eta$ and $\tau_{k}$ for conjugacy, respectively. The hyperparameters in the weakly informative inverse-Gamma prior distributions are set as $a_{0}=2$ and $b_{0}=1$ in our numerical studies. For the VB implementation, the variational distributions for $(\sigma^{(k)}_{y})^{2}$ in (3) are assumed to be inverse-Gamma distributions $\operatorname{I\Gamma}(a^{(k)},b^{(k)})$ for $k=0,\ldots,K$.
Since the assigned prior distributions are conditionally conjugate in the linear regression scenario, coordinate ascent variational inference (CAVI) (bishop2006pattern) can be used to solve (4), which iteratively optimizes each factorized term while keeping the others fixed.
@@ 3.2 Example: Logistic Regression
Unlike linear regression models, the logistic regression models do not enjoy direct conjugacy between likelihood and spike-and-slab priors, which makes MCMC and VB schemes intractable. One way to overcome this difficulty is to adopt Pólya-Gamma (PG) data augmentation (Polson2013). Specifically, we introduce a latent PG variable $\omega^{(k)}_{i}$ for each binary response $y^{(k)}_{i}$ and therefore the augmented likelihood $L({\cal D}\mid\omega)$ is given by
@@
where $\omega=(\omega^{(0)}_{1},\ldots,\omega^{(K)}_{n_{K}})$ with each $\omega^{(k)}_{i}$ following the PG distribution $\operatorname{PG}(1,0)$ for $i=1,\ldots,n_{k}$ and $k=0,\ldots,K$. The corresponding variational distributions of $\omega^{(k)}_{i}$’s are assumed to be $\operatorname{PG}(1,c^{(k)}_{i})$ with $c^{(k)}_{i}$ being the variational parameter. Data augmentation guarantees conjugacy so that CAVI can be utilized to solve (4).
@@ 4 Theoretical Results
In this section, we study the theoretical properties of CONCERT under the linear regression setting. We first provide theoretical results for the true posterior in Section 4.1 and further establish theoretical guarantees for the VB posterior in Section 4.2.
@@
Without loss of generality, we assume that $\sigma_{y}^{(0)}=\cdots=\sigma_{y}^{(K)}$ is known in the theoretical analysis. We consider the models with $|S|\leq L_{s}$ and $|T_{k}|\leq L_{t}$ for some large numbers $L_{s}$ and $L_{t}$. We first introduce some notations. Let $K]=\{1,\ldots,K\}$ and $N=\sum_{k=0}^{K}n_{k}$. Let $\boldsymbol{\theta}^{*}=(\boldsymbol{\beta}^{(0)*},\{\boldsymbol{\delta}^{(k)*}\}_{k=1}^{K})$ be the true parameters. We define the true signal set as $S^{*}=\{j\in[p]:|\beta_{j}^{(0)*}|\geq\zeta_{j}\}$ and the true non-transferable index set on source $k$ as $T_{k}^{*}=\{j\in[p]:|\delta_{j}^{(k)*}|\geq h_{j}^{(k)}\}$, where $\zeta_{j}$ is the signal level and $h_{j}^{(k)}$ is the heterogeneity level specified later. We use $s^{*}=|S^{*}|$ and $t_{k}^{*}=|T_{k}^{*}|$ to denote the true model size for the target and the true non-transferable model size of source $k$ for $k\in[K]$.
@@
For a matrix $A$ and an index set $D$, let the sub-matrix of $A$ with columns in $D$ as $A_{D}$. Let $I_{d}$ be the identity matrix of order $d$. We denote the complement of a set $A$ by $A^{\mathrm{c}}$. Denote the design matrix on the $k$th dataset as $X^{(k)}$. Then $\tilde{P}_{T_{k}}=X^{(k)}_{T_{k}}(X^{(k)\top}_{T_{k}}X^{(k)}_{T_{k}})^{-1}X^{(k)\top}_{T_{k}}$ and $\tilde{Q}_{T_{k}}=I_{n_{k}}-\tilde{P}_{T_{k}}$ are the projection matrices on the column space of $X_{T_{k}}^{(k)}$ and its complementary space, respectively. Let $\tilde{Q}_{T_{0}}=I_{n_{0}}$ and $\tilde{Q}$ be the diagonal block matrix of $\{\tilde{Q}_{T_{k}}\}_{k=0}^{K}$. We define the transformed design matrix as $\tilde{X}^{(k)}(T_{k})=\tilde{Q}_{T_{k}}^{1/2}X^{(k)}$ and the stacked transformed design matrix as $\tilde{\boldsymbol{X}}(T)=({X}^{(0)\top},\tilde{X}^{(1)}(T_{1})^{\top},\ldots,\tilde{X}^{(K)}(T_{K})^{\top})^{\top}\in\mathbb{R}^{N\times p}$. We further define $\Phi_{S,T}=\tilde{\boldsymbol{X}}_{S}(T)(\tilde{\boldsymbol{X}}_{S}(T)^{\top}\tilde{\boldsymbol{X}}_{S}(T))^{-1}\tilde{\boldsymbol{X}}_{S}(T)^{\top}$.
@@
Given a model $\{S,T\}$, we define the covariate-$j$-specific transferable set as $H^{(j)}_{S,T}=\{k\in[K]:j\in S\setminus T_{k}\}$. Denote $H_{S,T}=\cup_{j\in S}H_{S,T}^{(j)}$ as the index set of all least-transferable datasets given $\{S,T\}$, including all sources having at least one covariate similar to the target. We further define $n^{(j)}_{S,T}=n_{0}+\sum_{k\in H^{(j)}_{S,T}}n_{k}$ to be the useful sample size for covariate $j$ and $n_{S,T}=n_{0}+\sum_{k\in H_{S,T}}n_{k}$ as the sample size of all least-informative datasets given model $\{S,T\}$. Similarly to $\tilde{\boldsymbol{X}}(T)$ defined above, we use $\tilde{\boldsymbol{X}}_{H_{S,T}}(T)$ to denote the stacked transformed design matrix consisting of all least-informative datasets in $H_{S,T}\cup\{0\}$.
4.1 True Posterior Contraction
4.1.1 Oracle theoretical results with known $T^{*}$
@@
For some large constants $C_{1},C_{2}>0$ and all $k\in[K]\cup\{0\}$, it holds that (i) $\|{{X}}^{(k)}\boldsymbol{\beta}^{(k)*}\|_{2}^{2}\leq C_{1}n_{k}(\sigma_{y}^{(0)})^{2}\log p$; (ii) $\|\tilde{\boldsymbol{X}}_{S^{*c}}\boldsymbol{\beta}_{S^{*c}}^{(0)*}\|_{2}^{2}$$\leq C_{2}/2(\sigma_{y}^{(0)})^{2}\log p$, (iii) $\sum_{k=1}^{K}\|\tilde{X}_{T_{k}^{*\mathrm{c}}}^{(k)}\boldsymbol{\delta}_{T_{k}^{*\mathrm{c}}}^{(k)*}\|_{2}^{2}\leq C_{2}/2(\sigma_{y}^{(0)})^{2}\log p$; (iv) $\min_{j\in[p]}n^{(j)}_{[p],T^{*}}\geq c_{0}N$ for some $c_{0}\in(1/2,1)$.
Assumptions 1(i)-(ii) require that the order of the total signal strength of the $k$th dataset is no more than $n_{k}\log p$ and the non-influential signal strength in the transformed dataset is not large (yang2016on). Assumption 1(iii) imposes constraints on the magnitude of deviation within the transferable index sets across all sources. Notably, this assumption is less restrictive than the exact-transfer condition $\boldsymbol{\delta}^{(k)*}_{T_{k}^{*\mathrm{c}}}=0$, as it permits the incorporation of information from approximately similar sources without requiring perfect transferability. Assumption 1(iv) requires that the informative sample size along each dimension is comparable to the total sample size, excluding true models where only a small portion of sources are useful along each dimension.
@@
(ii) (Restricted eigenvalue) The gram matrix of the considered model space satisfies that $\min_{\{S:|S|\leq L_{s}\}}\lambda_{\min}({\tilde{\boldsymbol{X}}_{H_{S},S}^{\top}\tilde{\boldsymbol{X}}_{H_{S},S}}/{n_{S}})\geq\lambda_{l}$, $\lambda_{\operatorname{min}}(X^{(0)\top}_{S^{*}}X^{(0)}_{S^{*}})\geq n_{0}\lambda_{l}$, and for all $k\in[K]$ with $|T_{k}^{*}|\geq 1$, $\lambda_{\operatorname{min}}(X^{(k)\top}_{T_{k}^{*}}X^{(k)}_{T_{k}^{*}})\geq n_{k}\lambda_{l}$, for some $\lambda_{l}\in(0,1)$;
@@
(iv) (Weak correlation) For all $k\in[K]$, the transformed covariates satisfy that $\|\tilde{X}_{j}^{(k)}\|^{2}_{2}\geq d_{1}n_{k}$ (for $j\in T_{k}^{*\mathrm{c}}$) and $\tilde{\boldsymbol{X}}_{j}^{\top}(I-\Phi_{S})\tilde{\boldsymbol{X}}_{j}\geq d_{1}d_{2}n^{(j)}_{S}$ (for $j\in S^{\mathrm{c}}$) for all $S$ satisfying $|S|\leq L_{s}$, with some $d_{1},d_{2}\in(0,1)$.
Assumption 2 imposes conditions on the projected design, which are parallel to the assumptions for a single-task high-dimensional regression in the Bayesian literature (narisetty2014bayesian; yang2016on). Assumption 2(ii) allows the model size $S$ to grow slowly compared to $p$ and is satisfied when $L_{s}\leq\min\{p,n/\log p\}$. Assumption 2(iii) is mild and always holds under $C_{3}=O(\lambda_{l}^{-1}L_{s})$. Assumption 2(iv) imposes a weak dependence condition, similar to donoho2005stable and Castillo15. Under the special case of an orthogonal design, Assumption 2(iv) holds for $d_{1}=d_{2}=1$.
@@
The prior satisfies (i) $q_{0}\sim p^{-1}$; (ii) $n_{S^{*}}\eta^{2}\sim(n_{S^{*}}\vee p^{2r})$ for some $r>1+3/2C_{1}\lambda_{1}^{-1}+C_{2}+C_{3}$.
Assumption 3 is on the choice of prior hyperparameters, which is similarly assumed in narisetty2014bayesian and Narisetty2019ss.
@@
$\zeta_{j}^{2}\geq C_{\beta}\frac{(\sigma_{y}^{(0)})^{2}\log p}{n^{(j)}_{S^{*},T^{*}}d_{1}d_{2}},$
for $j\in S^{*}$, with constant $C_{\beta}=8({3}C_{1}{\lambda_{l}}^{-1}/2+r+3+(C_{2}+C_{3})/4)$.
Assumption 4 is parallel to the $\beta$-min condition in the high-dimensional literature for variable selection consistency (buhlmann2011statistics), which requires the signal strength on the significant index set to have a large enough magnitude. Note that the minimal signal strength restriction here is covariate-specific, depending on how many informative datasets there are along with the specific dimension.
@@
Suppose $\min_{\{S:1\leq|S|\leq L_{s}\}}n_{S}\gtrsim\log p$. Under Assumptions 1- 4, the posterior probability of the true model $S^{*}$ satisfies that,